1. Given π‘₯ = 60; and 𝑠 = 6, find the z-score that corresponds to each of the following

scores up to two decimal places.

a. π‘₯ = 70

b. π‘₯ = 58

2. Given πœ‡ = 72; and 𝜎 = 8, find the z-score that corresponds to each of the following

scores up to two decimal places.

a. π‘₯ = 68

b. π‘₯ = 80

3. Alex scored 90 during the first periodic exam in Mathematics and 88 during the

second periodic exam. The scores in first periodic exam have a mean πœ‡ = 83 and

a standard deviation 𝜎 =9. Scores in the second periodic exam have a mean πœ‡ =

80 and a standard deviation 𝜎 = 8. In which periodic exam was his standing better,

assuming that the scores in his periodic exams are normally distributed?

4. On a final examination in Biology, the mean was 75 and the standard deviation

was 12. Determine the standard score of a student who received a score of 60

assuming that the scores are normally distributed.

5. Given: πœ‡ = 64,𝜎 = 7. What is the raw score when 𝑧 = βˆ’0.76?



1
Expert's answer
2021-03-24T13:41:31-0400

1.

a.

"z=\dfrac{x-\bar{x}}{s}=\dfrac{70-60}{6}=\dfrac{5}{3}\approx1.67"

b.

"z=\dfrac{x-\bar{x}}{s}=\dfrac{58-60}{6}=-\dfrac{2}{3}\approx-0.67"

2.

a.

"z=\dfrac{x-\mu}{\sigma}=\dfrac{68-72}{8}=-0.50"

b.

"z=\dfrac{x-\mu}{\sigma}=\dfrac{80-72}{8}=1.00"


3. To answer this question we need to find the z score for each exam.


"z_1=\dfrac{x_1-\mu_1}{\sigma_1}=\dfrac{90-83}{9}=\dfrac{7}{9}\approx0.7778"

"z_2=\dfrac{x_2-\mu_2}{\sigma_2}=\dfrac{88-80}{8}=1"

We see that the highest z-score is 1, which means that Alex was his standing better in the second periodic exam.


4.

"z=\dfrac{x-\mu}{\sigma}=\dfrac{60-75}{12}=-1.25"



5.


"z=\dfrac{x-\mu}{\sigma}=\dfrac{x-64}{7}=-0.76"




"x=64-0.76(7)=58.68"